OpenStudy (anonymous):

Hello im new to open study an i need help with 3 qestions and its pretty urgent so if any one can help me i would be soo happy

OpenStudy (anonymous):

I will be happy to help :)

OpenStudy (anonymous):

WELCOME TO OPEN STUDY

OpenStudy (anonymous):

thanks by the way i just want help not the answers

OpenStudy (anonymous):

yeah of course

OpenStudy (anonymous):

anyway hers the fir question also thanks for helping me by the way

OpenStudy (anonymous):

An equation is shown below: 3(2x – 7) = 3 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation? Show your work. (6 points)

OpenStudy (anonymous):

you can solve it I will just guide you. :)

OpenStudy (anonymous):

im not sure how to do the problem do you think you can help me get on the right path because i have no ideo what to do

OpenStudy (anonymous):

3(2x – 7) = 3 Let's solve your equation step-by-step. 6x-21=3 6x-21+21=3+21 6x=24 solve for x

OpenStudy (anonymous):

That will be the only solution

OpenStudy (anonymous):

ok so can you tell me what you just did so i can learn from it please aslo since you were here to help i faned you and gave you a medal

OpenStudy (anonymous):

thanks but you din't have to I will give you ur medal back

OpenStudy (anonymous):

u are new

OpenStudy (anonymous):

oh ok so i geuse just gove it to you wehn were done

OpenStudy (anonymous):

don't givee me a medal I don't need it you are new

OpenStudy (anonymous):

i dont realy no how these things work i literly just made account like 7 min ago

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

LOL yeah

OpenStudy (anonymous):

do you have another question?

OpenStudy (anonymous):

uh yeah befor i go on to my next question you said solve x so what do i do add subtract multiply?

OpenStudy (anonymous):

24/6=x

OpenStudy (anonymous):

i realy suck in math its like my worst subject

OpenStudy (anonymous):

oh math is my set subject history is my worst

OpenStudy (anonymous):

so divide right

OpenStudy (anonymous):

best I mean

OpenStudy (anonymous):

yes divide

OpenStudy (anonymous):

realy historys my best lol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

next question?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

An equation is shown below: 9(3x – 16) + 15 = 6x – 24 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

while your solving that ill solve x

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

bythe way is it 6 divided by 24 or 24 divied by 6

OpenStudy (anonymous):

for x

OpenStudy (anonymous):

24/6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok i wont interfer anymore sorry

OpenStudy (anonymous):

Step by step solution : Step 1 :Simplify 27x-129 - 6x-24 Pulling out like terms : 1.1 Pull out like factors : 21x - 105 = 21 • (x - 5) Equation at the end of step 2 : 21 • (x - 5) = 0 Step 3 :Solve 21•(x-5) = 0 Solve : x-5 = 0 Add five to both sides and x-5 cancels so the solution is: 0+5=x

OpenStudy (anonymous):

i dont know what to do

OpenStudy (anonymous):

so 5+0= 5 so x=5

OpenStudy (anonymous):

soory to solve for u

OpenStudy (anonymous):

sorry I mean

OpenStudy (anonymous):

no i dont even know what you did im lost

OpenStudy (anonymous):

im realy realy realy bad at math

OpenStudy (anonymous):

oh with step number one I multiplied 9*3x and got 27x and I multiplied 9 and 16 and got 129

OpenStudy (anonymous):

I then pulled down the subtraction sign and got: 27x-129

OpenStudy (anonymous):

ok i think im starting to understand

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

pleasesssss

OpenStudy (anonymous):

then you add -24 to both sides: 27+-129+-6x+-24

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

combine like terms: (27x+-6x)+(-129+-24)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

also thank you and sorry for being so complicated

OpenStudy (anonymous):

you get 21x

OpenStudy (anonymous):

ok and then is that it

OpenStudy (anonymous):

21X*-5=0

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

0+5=x

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

do you kinda understand?

OpenStudy (anonymous):

ohhhhh yeah i get it now

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

it took me awile lol

OpenStudy (anonymous):

its okay at least you got it :)

OpenStudy (anonymous):

at first i was like what and then i was like ohhh now i get it

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

so was that for part 1 of question 2

OpenStudy (anonymous):

yes the explanation was x=5 is part 2

OpenStudy (anonymous):

oh ok so now the finle qestion yay i though i going to bomb this but because of you i have hope thank so much "cying on the inside"

OpenStudy (anonymous):

LOL your welcome happy to help :)

OpenStudy (anonymous):

"crying tears of joy"

OpenStudy (anonymous):

i dont know how to repay you man im just so happy but dont worry these are many tears "i told myself i wasnt going to cry but its just so beautiful"

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

da du da FINAL QUESTION "waring incoming major bossss" :boss Battle activated:

OpenStudy (anonymous):

An equation is shown below: 5x + 3(x – 5) = 6x + 2x – 15 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation. (4 points)

OpenStudy (anonymous):

here it is the finale question

OpenStudy (anonymous):

yayayayayayayayaya im almost done and then i can go to sleeep

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Reorder the terms: -15 + 5x + 3x = 6x + 2x + -15 Combine like terms: 5x + 3x = 8x -15 + 8x = 6x + 2x + -15 Reorder the terms: -15 + 8x = -15 + 6x + 2x Combine like terms: 6x + 2x = 8x -15 + 8x = -15 + 8x Add '15' to each side of the equation. -15 + 15 + 8x = -15 + 15 + 8x Combine like terms: -15 + 15 = 0 0 + 8x = -15 + 15 + 8x 8x = -15 + 15 + 8x Combine like terms: -15 + 15 = 0 8x = 0 + 8x 8x = 8x Add '-8x' to each side of the equation. 8x + -8x = 8x + -8x Combine like terms: 8x + -8x = 0 0 = 8x + -8x Combine like terms: 8x + -8x = 0 0 = 0

OpenStudy (anonymous):

remember im not as smart as you so nice slow OH MY GOOOOOOD IT SO COMPLEX

OpenStudy (anonymous):

Simplifying 5x + 3(x + -5) = 6x + 2x + -15 Reorder the terms: 5x + 3(-5 + x) = 6x + 2x + -15 5x + (-5 * 3 + x * 3) = 6x + 2x + -15 5x + (-15 + 3x) = 6x + 2x + -15 DO THIS BEFORE OTHER STEPS THAT I JUST POSTED

OpenStudy (anonymous):

my brain just stop working for a second there

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

DO THIS BEFORE OTHER STEPS THAT I JUST POSTED Simplifying 5x + 3(x + -5) = 6x + 2x + -15 Reorder the terms: 5x + 3(-5 + x) = 6x + 2x + -15 5x + (-5 * 3 + x * 3) = 6x + 2x + -15 5x + (-15 + 3x) = 6x + 2x + -15

OpenStudy (anonymous):

ok done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Are the steps kinda helping

OpenStudy (anonymous):

man your realy good at math

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

oh yes these steps are helping alot there still hard but i can finaly kinda understand and its not like going up to my teachers board and hes like solve this 30 secs and im like "what?"

OpenStudy (anonymous):

LOL